OpenStudy (anonymous):

If 0*2=0, then why is not 0/0 = 2???

OpenStudy (sierra!):

ummm That im not really sure with but if you give me a couple of minutes i can try to figure it out

OpenStudy (sierra!):

Yes im pretty sure it does not equal to 2

OpenStudy (anonymous):

but the transformation is fully legal, why should not the answer be right?

OpenStudy (sierra!):

Because i did work for the question and i did not get 2

OpenStudy (sierra!):

Therefore the question makes no sence

OpenStudy (sierra!):

Does any one agree with me?

OpenStudy (anonymous):

Well, it is a legal transformation of the statement, but dividing by zero contradicts how mathematics works.

Parth (parthkohli):

0*3 = 0 too. So is 0*4. So is 0*5.

Parth (parthkohli):

0/0 can be any number then, can't it?

OpenStudy (anonymous):

But as you divide 0 by 0 you have no value to distribute over 'infinity'

Parth (parthkohli):

Yeah, that's the reason why 0/0 is indeterminate.

OpenStudy (ikram002p):

ok this is very interesting and smart question, and itas answer depends on defining a group to work on. you can search for groups theory and abstract algebra for more info but i'll write why this is invalid operation:- lets assume we have an multiplication operation on Real then according to grouping axioms 0*r=0 and for any r in R-{0} there exist unique r^-1 =1/r such that r.r^-1=1 while 1 is the identity so according to this definition 1/0 is undefined value, they could not be 0*0^-1 =1 that satisfies ..... this is only a define and axiom matter its not about calculation but if we changes group Z_n for example we might make since of values 0^-1 but it wont be 1/0 it would be something else according to what n we have

Parth (parthkohli):

ok

OpenStudy (igreen):

Anything divided by itself is 1..

Parth (parthkohli):

no

OpenStudy (igreen):

Well, with whole numbers at least.

OpenStudy (igreen):

Same with 0

Parth (parthkohli):

no

OpenStudy (igreen):

Btw, why is your smartscore 20?

Parth (parthkohli):

i have called 911. you will be arrested soon.

OpenStudy (anonymous):

0/0 is not 1. It's an exception.

Parth (parthkohli):

and i'm not a smartscore obsessed noob.

OpenStudy (igreen):

Okay, curse Google

OpenStudy (igreen):

Just like how they say -1^2 = -1

Parth (parthkohli):

-1^2 is -1.

OpenStudy (anonymous):

(-1)^2 is 1, but -1^2 is -1

OpenStudy (sierra!):

And just saying for the record your smartscore means nothing you can be Albert Instine with a smartscore of 20

OpenStudy (sierra!):

It just means you dont help people out alot

Parth (parthkohli):

I'm completely sure that Albert Einstein helped people more than 99 SS though...

OpenStudy (sierra!):

Im just making a point

OpenStudy (ikram002p):

smart score means nothing :)

OpenStudy (sierra!):

Thank you!

OpenStudy (mathmath333):

http://en.wikipedia.org/wiki/Indeterminate_form

OpenStudy (ikram002p):

ugh :\ this is pure math lol calculus can't give explanation for such things it only used it like in limits why they call 0/0 indeterminate ?